OpenStudy (anonymous):

Solve one-fourth f plus three-fifths equals negative one-half for f. f equals negative eleven-fortieths f equals one-fortieth f equals two-fifths f equals negative four and two-fifths

OpenStudy (anonymous):

\[\frac{ 1 }{ 4 }f+\frac{ 3 }{ 5 }=\frac{ 1 }{ 2 }\]

OpenStudy (anonymous):

first subtract 3/5 from 1/2

OpenStudy (anonymous):

should be 1/10?

OpenStudy (anonymous):

but u have to turn 3/5 to 6/10 and 5/10

OpenStudy (anonymous):

*1/2 to 5/10... 5/10-6/10 is -1/10

OpenStudy (anonymous):

mmmm okay i get it.

OpenStudy (anonymous):

then divide 1/4 from -1/10

OpenStudy (anonymous):

dont get that part.

OpenStudy (anonymous):

hold on...leme explain...it will take a minute

OpenStudy (anonymous):

wouldnt it be 2 1/2?

OpenStudy (anonymous):

its -2 1/2

OpenStudy (anonymous):

o

OpenStudy (anonymous):

...but isnt one of the answers...

OpenStudy (anonymous):

but what would next step be.

OpenStudy (anonymous):

yep its not lol.

OpenStudy (anonymous):

it shouldnt be negative

OpenStudy (anonymous):

its positive i think.

OpenStudy (anonymous):

cause negative and positive make positive.

OpenStudy (anonymous):

no matter what thats what my teacher taught me.

OpenStudy (anonymous):

leme get a calculator up lol

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

its 2/5

OpenStudy (anonymous):

i divided wrong sorry lol

OpenStudy (anonymous):

2.5? u mean

OpenStudy (anonymous):

xdd

OpenStudy (anonymous):

all good.

OpenStudy (anonymous):

f equals two-fifths

OpenStudy (anonymous):

thanks :3

OpenStudy (anonymous):

ur welcome :x3

OpenStudy (anonymous):

mind helping a with like 3 questions?

OpenStudy (anonymous):

Renee took a long flight to visit her family that lives out of the country. She slept for the first quarter of the trip and watched movies for half of the remaining flight. If the total trip took nine and three-fourths hours, how long did Renee watch movies? three and twenty-one thirty-seconds hours three and one twenty-fourth hours three and five-twelfths hours five and two-thirds hours